OpenStudy (anonymous):

I need help solving my work. 12-a=11

OpenStudy (anonymous):

a = 1 12 - 1 = 11 :)

OpenStudy (anonymous):

i know its 1 but i need to show my work

OpenStudy (aum):

12-a=11 Add a to both sides: 12 = 11 + a subtract 11 from both sides: 1 = a

OpenStudy (anonymous):

can you write it out please

OpenStudy (aum):

12-a=11 Add a to both sides: 12-a+a = 11 + a 12 = 11 + a Subtract 11 from both sides: 12-11 = 11 + a - 11 1 = a or a = 1

OpenStudy (anonymous):

can you just write out the work with no explanation

OpenStudy (aum):

12-a=11 12 = 11 + a 12-11 = a a = 1